### v1

# What’s DAT? Three Case Studies of Measuring Software Development Productivity at Meta With Diff Authoring Time

Abstract.
This paper introduces Diff Authoring Time (DAT), a powerful, yet conceptually simple approach to measuring software development productivity that enables rigorous experimentation. DAT is a time-based metric, which assess how long engineers take to develop changes, using a privacy-aware telemetry system integrated with version control, the IDE, and the OS. We validate DAT through observational studies, surveys, visualizations, and descriptive statistics. At Meta, DAT has powered experiments and case studies on more than 20 projects. Here, we highlight (1) an experiment on introducing mock types (a 14% DAT improvement), (2) the development of automatic memoization in the React compiler (33% improvement), and (3) an estimate of thousands of DAT hours saved annually through code sharing ($>50$% improvement). DAT offers a precise, yet high-coverage measure for development productivity, aiding business decisions. It enhances development efficiency by aligning the internal development workflow with the experiment-driven culture of external product development. On the research front, DAT has enabled us to perform rigorous experimentation on long-standing software engineering questions such as “do types make development more efficient?”
Diff Authoring Time, DAT, Developer Productivity, Engineering Velocity, Continuous Experimentation (CE)
booktitle: Companion Proceedings of the 33rd ACM Symposium on the Foundations of Software Engineering (FSE ’25), June 23–27, 2025, Trondheim, Norwaycopyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM International Conference on the Foundations of Software Engineering (FSE); June 23rd-27th, 2025; Trondheim, NGisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineeringccs: Software and its engineering Software development methods
1. Introduction
For nearly two decades, Continuous Experimentation (CE) has been integral to external-facing product development in Big Tech (Kohavi et al., 2009; Tang et al., 2010; Karrer et al., 2021): Engineers frequently launch A/B experiments to evaluate new features, ensuring measurements are free of interferences. With a large user base, even small effect sizes are detectable. In contrast, internal development of developer tools used by thousands of developers mostly still relies on basic usage numbers and qualitative feedback (Forsgren et al., 2024). This is puzzling given the high value of engineering time, a major operational expense. To improve internal development, a metrics-driven culture is needed. However, measuring productivity accurately is challenging (Jaspan and Sadowski, 2019; Murphy-Hill et al., 2019; Beller et al., 2020), as traditional metrics are often inaccurate and easily gameable.
In this paper, we propose a solution inspired by Big Tech’s experiment-driven product culture. We introduce Diff Authoring Time (DAT) as a key metric for assessing development, not developer, productivity. DAT measures the time engineers take to develop a change, using a privacy-aware telemetry system integrated with version control, the IDE, and the operating system. We validated DAT through observational studies, surveys, visualizations, and statistics. Out of more than 20 studies completed with DAT, we highlight three, demonstrating DAT’s impact on productivity:
  1. (1)
    Introduction of mock types in Hack led to a 14% DAT improvement, marking a first in linking language-level features to productivity gains.
  2. (2)
    Quantification of productivity gains from a new feature in the React Compiler, auto-memoization, resulting in 33% productivity wins.
  3. (3)
    Development of a method to estimate efficiency gains from code sharing, showing a $>50$% improvement across frameworks, saving thousands of hours annually.
2. Background
Figure 1. Example illustration of the most relevant DAT-covered and uncovered activities for hypothetical diffs D123 and D987, and a comparison with other related metrics Code Gestation Time (CGT) and Time Spent by Diff (TSD).
2.1. Continuous Experimentation
The phrase “Continuous Experimentation” (CE) refers to the practice of regularly running experiments to validate one’s hypotheses, often simultaneously. While this approach has become widespread in large, software-intense organizations and has defined much of the engineering work in Big Tech (Kohavi et al., 2009), it is not ubiquitous (Mäntylä et al., 2022). Challenges include a lack of resources and knowledge on CE, particularly in startups.
Melegati et al. establish a “CE maturity model” with five stages of adoption and implementation (Melegati et al., 2020). Before this paper, Infra was at a similar stage to a typical startup: while Meta as a whole was naturally aware of CE (stage 1), it required a cultural mind shift to embrace it for Infra (stage 2). An organic need driven by lower observed organizational output combined with a top-down mandate initiated this. To instill the new mindset and work processes throughout the organization, a particularly rigorous validation (cf. Section 4) was necessary (stage 3). We derived results by focusing on a generally accepted measure (DAT), and integrating it into our already validated experiment infrastructure for product (stage 4). Highly skilled data scientists performed much of this work, leading to fast productionization of these results, including publishing the achieved wins in this paper (stage 5+).
2.2. Big Tech and organizational structure
By “Big Tech”, we refer to major software companies such as Alphabet, Apple, Microsoft, and Meta, with billions of users (Fernandez et al., 2020). Their internal development teams often support tens of thousands of engineers (Hedlund, 2021), optimizing tools and workflows to ultimately boost company-wide productivity.
While experimentation only makes sense given a large enough developer base, we believe DAT can apply to a multitude of stakeholders outside of Big Tech. Power analysis can reveal that sometimes a few dozen developers could be sufficient to detect effect sizes of the magnitude we present in this paper (Fitzner and Heckinger, 2010). It is also not strictly necessary to have full control over the complete workflow, as one is typically interested in one specific aspect when running experiments. Many consultancies, tool vendors, open source projects, and even some researchers (Beller et al., 2017) attract equally large or larger developer user bases, with the Eclipse IDE for Java surpassing 12 million monthly downloads in 2023, for example (Krill, 2023). Moreover, even in the absence of so many users, can the mental model behind DAT help shape how we as developers engineer with a data-driven mindset.
2.3. Tools and Processes at Meta
A diff represents a single code change including a natural text title, summary, and test plan (Beller et al., 2023), similar to a pull request (Gousios et al., 2014). Almost all code-based changes end up as (ideally, self-contained) diffs. A typical way of creating code diffs is by creating commits in an IDE such as VS Code (Marcey, 2019), using Sapling, Meta’s integrated, platform-independent version control client for multiple version control systems (Goode and Agarwal, 2014; Durham Goode, 2022). However, there are a number of other modalities for creating diffs, including bespoke tools (Murali et al., 2024). Code review of all diffs happens on a custom version of the open-source Phabricator tool (Kudrjavets et al., 2022; Beller et al., 2023). Overall, these characteristics make diffs a natural and high-coverage measurement unit for productivity and experimentation purposes.
Many productivity experiments make sense on the diff level, meaning that one developer can appear both in the A and B group, naturally removing one of the biggest confounders. For example, a diff-level experiment could be to test for the DAT effects of a new Continuous Integration (CI) signal, which would not affect the developer’s core workflow or state of mind if absent in the future, as one among many hundreds of CI signals. Similarly, switching out the base language model from LLaMA-33B to LLaMA-65B (Touvron et al., 2023) would likely be an opaque operation to a developer. Other experiments, however, (additionally) need to be on the cohort level—it makes little sense and is practically impossible to switch one developer back-and forth between two VS Code versions with completely different user interfaces, for example. A diff-based measurement is still essential in these cases, as the diff level provides normalization for many other confounders such as length, complexity, language, framework, or testing.
2.4. Related Metrics
Productivity metrics in Software Engineering have a long history (Brooks, 1974), with various metrics proposed, such as lines of code (Jones and Nelson, 1976), change count (Meli, 2001), and perceived productivity (Meyer et al., 2017). However, these metrics are often flawed, gameable, subjective, impossible to obtain accurately, or some combination of the above (Kitchenham, 1997; Petersen, 2011; Symons, 1988; Huijgens and Solingen, 2014). Some even suggest abandoning productivity measurement altogether (Ko, 2019).
Productivity frameworks such as SPACE (Forsgren et al., 2021) and DX Core4 (Noda et al., 2024) have recently provided frameworks on how to think about software development productivity at large, deferring the exact metric choice to their users. The metrics that are then often used tend to focus on easily measurable aspects of the development process, rather than the complexity of the authoring. For example, output metrics (such as lines of code, change count, …) emphasize the result of development while failing to capture the effort and complexity involved in the authoring lifecycle. Similarly, review time metrics, while good for measuring collaboration and delayed reviews, are not a measure of authoring time.
Therefore, there was a need to develop and test a metric that captured the complexity of authoring: DAT. DAT does this by measuring the active time taken to develop a diff while enabling insight into the components that impact it, such as IDE features and programming languages. For instance, feature experimentation using DAT enables teams to assess the impact of specific improvements on developer velocity.
We propose DAT as a conceptually simple, yet hard to implement, maximally objective metric that is not merely theoretical, but of high practical value. DAT supersedes Time Spent by Diff (TSD) and Code Gestation Time (CGT), illustrated in Figure 1. TSD averages coding time in a given period by the number of diffs published in that period, leading to misattribution and lacking diff-level precision. CGT measured time from coding start to diff landing, mixing active engineering with release time, complicating interpretation and lacking validation.
We adopt a pragmatic approach, focusing on accurate time measurement, as “time spent coding” aligns with developers’ productivity perceptions (Beller et al., 2020) and can be measured at scale.
3. Diff Authoring Time (DAT)
3.1. High-level description and vision
Diff Authoring Time (DAT) aims to capture all active human developer time spent on a particular diff, split by the different roles involved (e.g., author and reviewer), as depicted in Figure 1. It aims to make this information available in a privacy-aware manner. The key characteristic of DAT is the accurate representation of time spent in the coding editor, an Integrated Development Environment (IDE). DAT builds upon this to cover all diff-related activities, tools, and roles. Naturally, time off the computer, such as sketching whiteboarding or meeting time will always be impossible or difficult to track.
While there is an ongoing scientific debate on humans’ ability to effectively multitask (Koch et al., 2018; Skaugset et al., 2016), active human work on a computer tends to be highly sequential, even if such individual sequences can be extremely short (e.g., rapid application switching often happens in $<1s$): there can only be one mouse pointer, one cursor, and one application in focus at a given time, and such activity is normally related to one diff only.
This gives DAT some advantageous properties that lend themselves to validation: it is non-overlapping (per user), meaning that two diffs by the same author must not have any overlapping time. As a corollary, DAT cannot exceed 24 hours per day (though it is, in practice, much shorter), can sensibly be aggregated (as opposed to CGT), and the sum of DAT in one day can be bounded by TSD. In fact, the comparison with TSD can reveal (1) improvement potential for DAT’s matching algorithm or (2) potential for efficiency improvement, as coding-related activities did not result in productive output.
The goal for DAT is to be used purely as a productivity metric for tooling and workflows, not people. As a result, DAT and its underlying telemetry were designed with privacy in mind.
3.2. Core DAT algorithm: Precise matches
Figure 2. Explanation of the core DAT commit matching algorithm.
As the first two rows in Figure 1 depict, the core authoring of a diff starts with the initial trace of an activity that we can connect with this diff. Typically, this is the start of the first IDE session—most often in VS Code—in which the first Sapling commit is made that is then later connected to that diff. DAT ends with the final IDE session connected to the diff. DAT includes only the time actively working on this diff in-between the two timestamps. DAT is effectively a non-overlapping function of a diff, meaning that DAT(D123) $\cap$ DAT(D987) = $\emptyset$. We further split DAT into an author and reviewer time component. The core DAT algorithm prioritises precise, high-confidence matching—it omits IDE activity that does not result in a commit, or activity for commits that do not result in a diff. It only operates on tools that allow such high confidence attribution: IDEs and the Phabricator Diff review tool.
We obtain basic activity and foreground application through the privacy-aware OS-level telemetry. We then correlate it to commits using telemetry from the Sapling plugin. Sapling has the advantage of tracking the underlying repository, meaning that any interaction with it (e.g., also from the command line) is captured.
Figure 2 depicts the fundamental principle of DAT’s diff-to-session matching algorithm using the example of two diffs D123 and D987 that are developed sequentially in the same VS Code session: on top, the figure shows human activity as work sessions, along with point-in-time interactions with version control. It also shows the state of the repository (the output of sl status, similar to git status)—the current commit tracked by the version control system. The figure depicts the principle that any commit $CH_{x}$ is attributed on the time axis to the IDE session $s_{(x-1)}$ that precedes it. Thus, at its core, the DAT algorithm performs a 1-left-shift of the underlying commits with regard to the telemetry sessions. For example, the commit $CH1$ made at time $t2$ is attributed to the VS Code session $s1$ that happened from $t1$ to $t2.$ This is because any change developed starting from $t1$ would become part of the commit $CH1$. Similarly, both $s2$ and $s4$ are mapped to $CH2$. $s2$ is interrupted by the offline activity $s3$, which does not count toward DAT.
There are two corner cases: (1) Any time a developer opens VS Code, an initial automatic checkout happens ($CH0$ in Figure 2). As the developer does not modify this pre-existing commit, the commit event from this automatic checkout is left-shifted away by the DAT algorithm. Similarly, DAT removes the later manual checkout of the pre-existing commit $CH7$ via explicit filtering. The commit $CH8$, which becomes $CH7$’s successor, is equally attributed to D987. (2) For the “trailing session” $s7$, there is no commit mapping because any work done there would not actually be saved as a commit.
We now have a precise commit-to-session mapping (hence the algorithm name “precise matches”), but not yet a mapping to diffs. This, though, is trivial to achieve, as Version Control provides us with a list of all commits that make up a diff.
3.3. Additional DAT Heuristic: Anchor sessions
DAT includes additional heuristics beyond precise matches. These were developed to capture more of the time surrounding diff creation, especially in tools not supported by DAT telemetry. The idea generally is to map coding activity preceding a precise match to the same diff as the precise match, recognizing that related activities often occur sequentially.
For instance, in Figure 1, terminal work cannot be precisely matched due to lack of Sapling integration, but Anchor-DAT maps it to D123 because it precedes a precise match. This conservative method captures anchor sessions leading up to a diff, which are otherwise hard to trace. Activity is captured only for a limited time before a precise match and only if it involves coding-related tools. The definition of coding-related tool is shared with TSD, thus duration covered by DAT is a true subset of the duration covered by TSD.
4. Validation
4.1. User Experience Research Study
Our first step in validating DAT was to create a ground truth (GT) dataset of how long developers spent on a diff. This allowed us to assess our current DAT version’s accuracy by comparing it with GT data and observing trends as we improved the algorithms.
We conducted a User Experience Research (UXR) study, screen-recording volunteer developers working on a diff. To counteract biases from screen-recording and volunteering, we used a stratified random selection of developers and diffs for broad coverage. A UX researcher transcribed video actions, noting tool use and relevance to diff development. In the future, this could train an AI to derive ground truth timings from video data.
We analyzed over a dozen developers’ diffs, achieving an average accuracy of over 90%, meaning the computed DAT was on average within ±$\pm 5\%$ of its GT value. In later DAT versions, the algorithm’s accuracy often surpassed human labelling, as developers rapidly switched between diffs, challenging human annotators.
4.2. Large Scale Survey
Figure 3. 99th percentile winsorized average DAT for different algorithms, across a range of several months. The figure clearly shows the stability of DAT as a top-line metric.
Figure 4. Human DAT estimates (solid) and actual DAT values for an earlier DAT version (dashed).
While the UXR study was crucial for testing and development, it was limited in scale. Moreover, acceptance of a metric depends on how well it matches experience. Therefore, we conducted a survey asking developers about the time spent on a diff. This survey was integrated into the Phabricator tool, appearing once a diff was shipped to ensure responders had the full development context fresh in their minds.
We received responses for 968 unique diffs to compare with DAT values. As Figure 4 shows, DAT and survey data generally align, with developer-reported times being slightly larger, a trend also seen in the UXR study (Section 4.1). The difference between survey and DAT values increases for longer diffs, but this drift has been mitigated with newer DAT versions.
4.3. Descriptive Statistics
We performed two main validations for DAT at an aggregate level—diff coverage and diff duration. These validations were intended to verify that DAT covers the expected number of diffs and that the overall time spent in these diffs yields reasonable values. As a sanity check, the coverage and duration validations were done by comparing the current DAT version to its predecessor. The duration validations also included comparison between user-level DAT and TSD (cf. Section 2.4).
We performed several tests based on our expectations of how much additional heuristics should increase DAT. First, we expected anchor sessions durations to be higher than precise match durations. Second, we expected that the sum of all precise matches and anchor session durations never be larger than the sum of all TSD for a user during a particular time window. Any exceptions would only be possible if there is a mistake in the data or our algorithms, for example double-counting time. A final demand on any top-line metric is that it be stable over time—Figure 3 shows this.
Before releasing DAT, we ensured it passed all these checks. We found that DAT covers 87% of all eligible diffs. DAT’s duration has a right-skewed distribution, which means a small number of extreme values can significantly impact the mean. To account for slower diffs while still being sensitive to changes in the distribution, we use the 99th percentile winsorized mean to report on Anchor-DAT, which was very plausible with regard to other validations.
4.4. Time-Line Visualization
Figure 5. A sample timeline of activities (raw telemetry, bottom row), and different DAT algorithms.
Figure 5 visualizes an example diff’s DAT and how raw telemetry is covered by an additional heuristic, Anchor-DAT. DAT uses raw telemetry sessions, including start and stop times of coding tools, shown in the bottom row and maps each session to a diff, if possible. The raw sessions already include noise-removal, like pauses for inactivity (e.g., around 12:45). During DAT additional heuristic development, we also cross-validated with authors to ensure extrapolations were sensible. These checks helped refine our understanding, re-evaluate heuristic assumptions, and adjust parameters like time thresholds.
5. Experiment: Typed Mocking
1 // Existing test case
2 class AgeTest extends MetaTest {
3 public function testYoungCantVote(): void {
4 self::mock(Database::queryUserData<>, (\$_user) ==> dict["name" => "Alice", "age" => 16]);
5 expect(Application::canUserVote("Alice"))->toEqual(false);
6 }
7 }
8
9+class UserData {
10+ function __construct (public string \$name, public int \$age ) {}
11+}
12
13-class Database { function queryUserData(string \$user): dict<string, mixed>; }
14+class Database { function queryUserData(string \$user): UserData; }
15
16 class Application {
17 function canUserVote(string \$user): bool {
18 \$user_data = Database::queryUserData(\$user);
19- return (int)\$user_data[’age’] >= 18;
20+ return \$user_data->age >= 18;
21 }
22 }
(a) An exemplary code change (“diff”) introducing a new type, UserData.
// Testing framework, old, untyped mock function
class MetaTest {
function mock(mixed \$fn, mixed \$callable);
(b) The old mocking function leads to a TypehintViolation in AgeTest at runtime.
// Testing framework, new, typed, mock function
class MetaTest {
function mock<T>(FunctionRef<T> \$fn,
T \$callable): void;
(c) The new mocking function leads to a static type (lint) error in AgeTest at development time.
Figure 6. Comparison of untyped and typed mocking APIs.
5.1. Context
Hack is one of our server-side languages. Originating from PHP, it combines a JIT-based runtime (Ottoni, 2018) with a gradual, object-oriented type system (providing static correctness guarantees similar to TypeScript for JavaScript). Hack developers can depend on commonly shared frameworks, including mocking for testing (Spadini et al., 2017). While CI forces test execution for every diff, developers can additionally run them locally if they so wish.
Hack code is statically typed to varying degrees. In the presence of suitable type annotations, Hack has the ability to detect certain type errors through static analysis, which are reported to developers directly in VS Code. In addition to static analysis, the Hack runtime enforces the arguments passed to functions, and ensures that they match the declared parameter types, throwing a dynamic TypehintViolation-exception in case of a mismatch. In particular, the standard mocking framework used to expose an untyped API, using the catch-all type mixed in Listing 6(b). As a result, its use did not enjoy any static guarantees. Values flowing through ‘mixed‘ could encounter runtime enforcement boundaries, potentially leading to runtime crashes.
5.2. Problem Statement
In the example diff in Listing 6(a), a developer improves the Application to be more object-oriented, introducing a new UserData class, and altering existing code (lines 13 and 19) to make efficient use of it. However, the developer might not be aware of the existing test.
In the case of the old, untyped mocking framework (Listing 6(b)), they would learn that they had broken a test only after CI ran, introducing significant delays: since the mocked method-body now returns a dict value where UserData is expected, the Hack runtime throws a TypehintViolation. This leads to an additional cycle of checking out a diff, reproducing and debugging the unexpected failure, coming up with the right fix, and verifying it. The lack of static types for mocking has proven to be a hindrance to productivity, as indicated by the disproportionately high number of TypehintViolation-exceptions from CI logs.
To avoid this degraded experience, we built a typed variant of the existing mocking framework. Listing 6(c) shows the new typed mock function. The Hack type checker will immediately report a type mismatch as a type error in the developer’s IDE because the closure type is incompatible, without the need for the developer to know about or execute any potentially existing tests.
5.3. Experimental Setup
Our experiments showed that we would be able to automatically migrate the majority (82%) of calls into the untyped mocking API onto the new typed variant (migratable call sites). The remainder of call sites (or unmigratable call sites) exercised the mocking API in unconventional ways and could not be auto-migrated.
We proceeded to first migrate a random selection of half of the migratable call sites ($\sim$169,000 or 42% of all call sites in the codebase). During a period of three weeks, we categorised all committed diffs:
  • Control Group (53.1% of diffs): Diffs that modified at least one migrated, but no unmigrated file.
  • Test Group (24.5% of diffs): Diffs that modified at least one unmigrated, but no migrated files.
  • Mixed Group (22.4% of diffs): Diffs that modified at least one migrated and one unmigrated file.
To estimate the effect of typed mocking, we employ a two-tailed Welch’s t-test, testing on the mean DAT difference between control and test groups. We discarded the mixed group as it contained diffs that touched both migrated and unmigrated files. Because bigger diffs naturally require more coding time, we controlled for the diff size confounder using stratification.
5.4. Results
Table 1. Typed Mocking Experiment results for t-tests with diff size as stratum.
# of files changed % DAT saved $\frac{|Control|}{|Test|}$ $p$
1 $\sim 22\%$ $2.5$ $<0.001$
2 $\sim 22\%$ $2.3$ $<0.001$
3 $\sim 19\%$ $2.2$ $0.003$
4 $\sim 24\%$ $2.0$ $0.001$
Variable $\sim 14\%$ $2.2$ $<0.001$
Table 1 summarizes the results of the stratified t-tests, revealing that use of the typed mocking API significantly reduced the time to write diffs, even after controlling for the number of files in each diff. The table shows the percentage of DAT saved, the relative control and test group sizes, and the p-value for each diff size category; diff size is measured by the number of files changed in a diff.
For diff sizes up to 4 files, the typed mocking API resulted in an approximate 19–24% reduction in DAT, with all diff size categories showing statistically significant reductions. When including all diffs regardless of diff size, we measure a statistically significant reduction of approximately $14\%$ DAT ($p<0.001$). This is lower than the values in Table 1, as the inclusion of larger diffs (with more than four files changed), which only tangentially touch mocking-related files, dilutes the real savings seen in lower diff size strata.
To the best of our knowledge, this is the first time a controlled, large-scale experiment demonstrated the productivity benefits of typing in industry. While the results of this experiment have helped to quantify the direct effect of a Hack language feature and its use on the company’s productivity, the wider organizational decision-making processes also benefit. Moreover, this result demonstrates that investing in low-latency static analysis tooling to shift correctness signals early in the development cycle can lead to substantial improvements in developer productivity, especially in large code bases.
6. Case Study: React’s Auto-Memoization
6.1. Context
React is Meta’s JavaScript framework for developing dynamic user interfaces (Gackenheimer, 2015). Since its open-source release in 2013, it has become a standard for web and mobile apps (Facebook Inc., 2013). React components generate user interfaces based on the application’s current state. When the state changes, React updates the interface by re-evaluating components. To optimize performance, developers must use memoization to prevent recomputing unchanged components.
6.2. Problem Statement
React developers can memoize in two ways:
  1. (1)
    Manual memoization, which involves writing code to save component states manually and for every component independently.
  2. (2)
    Auto-memoization, available in recent React versions, uses a new Meta-built compiler feature to automatically memoize components, representing a significant investment by the React team.
Manual memoization is thought to be time-consuming, prone to bugs, and does not scale, but the magintude of its impact on productivity remains unknown.
6.3. Methodology and Results
Figure 7. Auto-Memoization Case Study results show beanplots of DAT for manual (left) and automatic memoization (right) in a small subset of React diffs.
To understand what difference the new React compiler might make to development velocity, we first compared diffs with auto-memoization to those without. We find that the DAT median, quartiles, and skew are significantly higher for diffs with manual memoization.
To account for confounding variables such as lines of code, developer tenure, organization, and time (modelled as month of the year), we built a mixed effect regression model. We computed %$\Delta\mbox{DAT}$ to measure differences in mean DAT distributions, shown in Figure 7. For non-randomized data, we used the Wasserstein distance to measure the true difference between groups. This metric provides the minimal transport distance between two probability distributions (Wikipedia, 2024). It is robust to noise and outliers, making it ideal for comparing distributions with different shapes or scales. The final results put %$\Delta\mbox{DAT}$ between diffs with manual memoization and without memoization at 33%.
7. Case Study: Quantifying Efficiency Gains from Code Sharing
7.1. Context
Meta’s development frameworks enable cross-platform code sharing, allowing code to be written once and deployed across multiple apps and platforms, reducing redundant efforts. One well-known example of such a framework is the open-sourced React framework (Facebook Inc., 2013). Using such frameworks supposedly saves time compared to writing platform-specific code, because it only needs to be written once, but can be deployed to a number of different platforms. Developers report feeling more efficient with these frameworks, yet quantifying these gains before DAT was impossible.
7.2. Methodology
To quantify $DAT_{saved}$, we subtract actual DAT development time from an estimated counterfactual DAT without code sharing.
Counterfactual DAT aims to provide a realistic estimation for how much DAT diffs would have taken if code sharing was not used. While DAT for the code-shared diffs is easy to query, we cannot directly measure counterfactual development time without sharing because these hypothetical diffs were never written. Instead, we estimate it by identifying the platforms (such as Android) and apps (such as Instagram or Threads) targeted in each code-shared diff. To obtain a baseline for each platform-app combination, we compute a trimmed mean DAT for unshared diffs. We normalize by diff size, dividing diffs into terciles based on modified lines of code.
The total counterfactual DAT is then the sum of all targeted platform-app combinations in the code-shared diff, controlled for diff size. A consequence of this is that $DAT_{saved}$ could be negative, in case a code-shared diff incurs overhead possibly due to the complex nature of shared libraries. Conversely, the higher the platform-app count is, the greater the potential for actual savings.
7.3. Results
We initially applied this method to two major code-sharing frameworks, covering several thousand of code-sharing diffs and counterfactual diffs monthly. Our results indicate that in this case alone, we save thousands of hours of DAT annually, a relative improvement to non-shared development of over 50%. When we set out, it was unclear whether there were substantial savings from code sharing; DAT enabled a precise statistical comparison, moving beyond anecdotes and intuition.
8. Threats to Validity
As a top-line metric DAT strikes a balance between support of all engineering workflows at the company (external validity), and precision for each diff (internal validity). These objectives are often at odds, as the example of Anchor-DAT shows: extrapolation increases tool coverage at the expense of potentially covering activity not related to this diff.
To mitigate this, we performed a quadruple validation of DAT (see Section 4), including large-scale surveys and end-to-end tests with GT (construct validity). Moreover, we plan to increase accuracy by supporting more tools and IDEs in a native way. For instance, we have already seen the beneficial effects of the addition of Android Studio as a natively supported IDE.
9. Discussion
9.1. Going beyond velocity
This paper demonstrates the broad applicability of DAT in practice, enabling Software Engineering-related case studies and experiments on different levels, from language features and compiler improvements to a generalized method to calculate framework efficiency gains. While DAT offers insights into velocity, there are other important aspects to software development such as quality (Cavano and McCall, 1978). Our experiments and case studies have shown, though, that these often go together: a DAT improvement frequently co-occurs with other desired properties, such as more static safety guarantees through typed mocking in Section 5).
Our studies have established that language and framework improvements (e.g., the React compiler improvements in Section 6) provide large velocity improvements. However, for an end-to-end analysis, their initial development efforts need to be balanced against the later DAT savings, ensuring the benefits outweigh the initial added development overhead. Fortunately the initial development cost can be measured with DAT, too, providing a net benefit analysis of these features. Having performed over 20 DAT experiments and case studies, we are now in a position to inform planning and prioritization of development in a more rigorous, data-driven way.
9.2. Dealing with small effect sizes
Another important consideration is the challenge of measuring a statistically significant change in DAT, especially when working with small sample sizes. To overcome this, experiments often need to run for extended periods to account for variability in workflows and usage patterns. This distinguishes internal developer experimentation from external consumer experimentation, where changing the position of a button by a few pixels can yield measurably different results (Goodui.org, 2020). The fidelity of DAT might not lend itself to such types of features or improvements.
We believe that this could also be attributed to the fact that developers are highly cognitively engaged in their work and constantly striving to find optimizations themselves. Their optimization scope and creativity are much larger than the often confined ecosystems in the consumer-facing product world, which equally means that capturing them in a metric makes it so much harder. An example of this would be developers that continue to use their favorite bespoke editor, even if there is only one officially supported workflow to creating diffs, and copy their edits in for example. The DAT for such copy-and-paste might be extremely short, and thus be of no value for experimentation.
In such cases, engineering teams might show impact by improving an operational metric more dependent on their changes, for example, number of successful searches. While a statistically significant, positive change in this metric might not yield statistically significant effects on DAT, the team could rely on a previously demonstrated, principal correlation between that metric and DAT to claim productivity improvements.
9.3. Embracing an experiment-first culture
An experimentation workflow requires time and rigor, and may even require a shift in the behavior of Infrastructure engineers, who might not be used to this mindset. As an example, the experiment in Section 5 took into account the DAT experiment and involved a staged migration of mock call-sites, to ensure DAT would be correctly measured. This additional effort is often only worth the overhead if engineers expect a significant DAT reduction. This however assumes that a feature is known to be strictly beneficial a-priori, which is often only a post-hoc learning. To design code in a DAT-first way can significantly alter the feature itself, as was the case in Section 5, where we did not migrate all untyped mock call sites at once. In this sense, there could be an opportunity cost to running an experiment.
Finally, DAT can serve as an enabler for greater inter-team mobility. It aligns the experiment-driven consumer-facing product development workflow with how internal Infrastructure teams operate, thus making it easier to collaborate or switch between teams.
9.4. Avoiding metric-blindness
At the same time it is important to recognize that metric movement should not become a condition or even barrier to pursuing revolutionary ideas. Some improvements may not result in (measurable) DAT reductions, either because they cannot be measured or because no valid experiment can be setup. It remains paramount that we as software builders continue to listen to our expert opinion on future innovations as well as current workflow hindrances, next to assessing more large-scale sentiment and qualitative feedback from fellow developers. We must maintain a balance between leveraging DAT for data-driven decisions and fostering a culture where innovative ideas can thrive. Consider the example of coding agents that ingest a development task and work on it in a human-like way (Hassan et al., 2024). Due to their current speed limitations, they might initially have adverse impact on DAT (i.e., a developer can perform the same actions in a much shorter DAT-equivalent than the agent), though their long-term trajectory looks extremely promising.
Despite these challenges, DAT provides key insights into developer velocity, and is measured against when improving developer workflows. By precisely capturing the previously opaque authoring time, DAT enabled a paradigm shift for Infrastructure developers, bringing them into a more data-driven and objective world.
9.5. Experimenting at Industry-Scale
Randomized controlled trials and controlled experiments are often called the gold standard in science (Manzi, 2012). Unfortunately, they continue to be a rarity in Software Engineering (Sjøberg et al., 2005; Stol and Fitzgerald, 2018).
While a relatively large body of such studies exists in the space of types (Hanenberg, 2010; Endrikat et al., 2014; Kleinschmager et al., 2012), understandably, one limitation of many previous studies was the reliance on students as experiment subjects (Carver et al., 2004) and the inability to do like-for-like comparisons between control and experiment groups. This might be a reason why studies came to contradictory results on the effectiveness of types. Our experiment here suffers no such limitations and is, to the best of our knowledge, the first one to be conducted in a large-scale, state-of-the-art, industrial setting with real developers on real-world problems. Partially contradicting previous findings (Hanenberg, 2010), our results show strong productivity improvements when using types. It is possible that the nature of our experiment, namely the introduction of typed mocking in a large-scale system, is more comparable to previous studies which favored types as system or task complexity increased (Stuchlik and Hanenberg, 2011; Endrikat et al., 2014; Gao et al., 2017). While this hypothesis seems very plausible, we need more rigorous experiments in different settings to assess whether static types hinder ad-hoc, quick development, but are actually a productivity improvement in larger projects. It is also unclear where the threshold between these two extremes lies.
For code sharing and compiler improvements, there exists as body of largely anecdotal or post-hoc empirical analyses touting their benefits: These range from cost reduction to (Haefliger et al., 2008; Bhattacharya and Neamtiu, 2011; Mockus, 2007). However, these works are often somewhat older, seldom contain concrete efficiency gains, stem mostly from Open-Source Systems, or lack a focus on (time-based) developer productivity. Moreover, we are unaware of controlled experiments, making the studies in this paper possibly the first in the respective areas.
10. Future Work
We are continuously evolving DAT by increasing coverage horizontally and vertically. Horizontally, we will expand DAT beyond diffs to include other artifacts such as documents and tasks, creating a general time-based measurement framework for experimentation. Vertically, we aim to cover more tools and IDEs natively, having to rely less on heuristics.
As AI advances in software development (Khemka and Houck, 2024), we believe DAT will remain relevant. Coding agents learn from human interactions, and DAT could guide the optimization (or loss) function toward higher productivity.
11. Conclusion
We conceptualized DAT as a practical productivity metric, implemented it, and used it successfully in more than 20 projects across Meta, of which we highlight three in this paper:
  1. (1)
    Mock types in Hack, denoting a 14% DAT improvement; to the best of our knowledge, this was the first, large-scale, real-world experiment accounting for relevant confounders that demonstrated productivity improvements by static typing;
  2. (2)
    Auto-memoization in the React compiler, showing a 33% DAT improvement;
  3. (3)
    Increased use of code sharing, showing a $>50\%$ DAT improvement, saving thousands of hours annually.
In conclusion, DAT has enabled Meta’s Infrastructure organization to adopt a metric-first development approach, aligning it closer with product teams. This way, DAT enabled a paradigm shift in how Infrastructure engineers work, report, and plan. DAT also paved the way toward rigorous experimentation on long-standing software engineering debates, such as “do types make development more efficient,” marking significant business and research advancements.